Find the equation of the ellipse whose foci are at(0,-5) and (0,5) and the length of its major axis is 14


1
Expert's answer
2021-02-03T00:15:24-0500

Solution:

Given, foci are (0, -5) and (0, 5).

Here, "c=5"

Also, length of major axis is 14.

So, "2a=14"

or, "a=7"

So, this ellipse lies along y-axis, i.e. major axis is y-axis.

Also, (0, 0) is the centre of the ellipse.

Consider its equation: "\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1" ...(i)

We know that, "b^2=a^2-c^2"

"\Rightarrow b^2=7^2-5^2" [Putting values]

"\Rightarrow b^2=49-25=24"

Putting values of a and "b^2" in (i), we get

"\dfrac{x^2}{7^2}+\dfrac{y^2}{24}=1"

"\Rightarrow \dfrac{x^2}{49}+\dfrac{y^2}{24}=1"

Answer: "\dfrac{x^2}{49}+\dfrac{y^2}{24}=1"